# 3D Bivariate Spatial Modelling of Argo Ocean Temperature and Salinity Profiles

### v3

@@ 1 Introduction
The core Argo data products are temperature and salinity–the two fundamental variables from which many other oceanographic variables can be derived, including freezing point, electrical conductivity, and viscosity \@BBOPcitep\@BAP\@BBN(pawlowicz2013key)\@BBCP, ocean heat content and potential density \@BBOPcitep\@BAP\@BBN(yarger2022functional)\@BBCP, and tropical cyclone-induced ocean thermal response \@BBOPcitep\@BAP\@BBN(hu2020spatio)\@BBCP. Temperature and salinity are central to understanding the oceans’ physical properties and dynamics: their distributions influence ocean circulation \@BBOPcitep\@BAP\@BBN(chen2022effects; gangopadhyay2022introduction)\@BBCP, shape climate processes \@BBOPcitep\@BAP\@BBN(olson2022effect)\@BBCP, and alter biogeochemistry \@BBOPcitep\@BAP\@BBN(galan2021argo; ding2022effects)\@BBCP. They also support diverse applications, including hydroacoustics \@BBOPcitep\@BAP\@BBN(jana2022sound)\@BBCP, offshore wind farm design \@BBOPcitep\@BAP\@BBN(escobar2016influence)\@BBCP, and assessments of environmental stressors in marine and freshwater systems \@BBOPcitep\@BAP\@BBN(walker2020warmer)\@BBCP.
@@ 1.1 Spatial covariance models in 2D and 3D space
$C_{ij}(\mathbf{s}_{1},\mathbf{s}_{2})=\rho_{ij}\sigma_{i}(\mathbf{s}_{1})\sigma_{j}(\mathbf{s}_{2})\mathcal{M}_{\nu_{ij}}\left\{(\mathbf{s}_{1}-\mathbf{s}_{2})^{\top}\mathbf{D}_{ij}(\mathbf{s}_{1},\mathbf{s}_{2})^{-1}(\mathbf{s}_{1}-\mathbf{s}_{2})\right\},$ (5)
where $\mathbf{D}_{ij}(\mathbf{s})=\frac{1}{2}\left\{\mathbf{D}_{i}(\mathbf{s}_{1})+\mathbf{D}_{j}(\mathbf{s}_{2})\right\}$ is a $d\times d$ positive definite kernel matrix that controls spatially varying anisotropy, for $i,j=1,\ldots,q$. Despite this innovation, extending the Matérn model to include nonstationary colocated correlations $\rho_{ij}(\mathbf{s})$ remains challenging. One proposed remedy is to replace the constants $\sigma_{i}^{2}$ and $\beta_{ij}$ with kernel-smoothed nonparametric estimators \@BBOPcitep\@BAP\@BBN(kleiber2012nonstationary)\@BBCP:
@@ 2.1 Spatial Interpolation of Argo Data
One widely used Argo product is the Roemmich–Gilson Argo climatology from the Scripps Institution of Oceanography \@BBOPcitep\@BAP\@BBN(roemmich2009)\@BBCP, which employs weighted local regression. Let $\mathbf{s}=(L,l,p)$ denote spatial location, where $L$ is latitude, $l$ is longitude, and $p$ is pressure (commonly treated as depth in the upper ocean). Let $t$ denote time in yeardays (days since January 1 of a given year). For a reference location $\mathbf{s}_{0}=(L_{0},l_{0},p_{0})^{\top}$ and reference time $t_{0}$, \@BBOPcite\@BAP\@BBNroemmich2009\@BBCP fit the mean function:
$\mu_{\mathbf{s}_{0},t_{0}}(\mathbf{s},t)=\beta_{0}+\beta_{1}(L-L_{0})+\beta_{2}(l-l_{0})+\beta_{3}(L-L_{0})^{2}+\beta_{4}(l-l_{0})^{2}+\beta_{5}(p-p_{0})\\ +\beta_{6}(p-p_{0})^{2}+\sum_{k=1}^{6}\gamma_{k}\sin\left(2\pi k\frac{t}{365.25}\right)+\sum_{k=1}^{6}\delta_{k}\cos\left(2\pi k\frac{t}{365.25}\right)$ (7)
using observations from the 100 nearest horizontal neighbors, pressure levels adjacent to $p_{0}$, and times within 12 months of $t_{0}$. Here, $\beta_{0}$ and $\beta_{k}$, $\gamma_{k}$, and $\delta_{k}$, $k=1,\ldots,6$, are scalar coefficients. The residuals were modeled using the spatial covariance function:
@@
Most studies discretize the vertical domain into fixed pressure levels and interpolate each layer independently, which requires preprocessing irregular profile data and can introduce systematic error \@BBOPcitep\@BAP\@BBN(kuusela2018locally; hu2020spatio)\@BBCP. To avoid this, \@BBOPcite\@BAP\@BBNyarger2022functional\@BBCP treated Argo profiles as functions of pressure and modeled the local mean as
@@ 3.1 Multivariate Nonstationary Spatial Models in 3D Domain
Furthermore, suppose the cross-covariance function $C_{ij}^{X}$ is chosen to be the parsimonious Matérn cross-covariance function in (1). Then, the multivariate process $\mathbf{Z}=(Z_{1},Z_{2},\ldots,Z_{q})^{\top}$ has a nonstationary cross-covariance function:
@@
$K_{ij}^{1}=\frac{\alpha_{ij}}{4}\sum_{k=1}^{K}\{a_{i,k}(L_{1})a_{j,k}(L_{2})h_{1}h_{2}-b_{i,k}(L_{1})b_{j,k}(L_{2})h_{3}^{2}-c_{i,k}(L_{1},p_{1})c_{j,k}(L_{2},p_{2})h_{4}^{2}\\ -a_{i,k}(L_{1})b_{j,k}(L_{2})h_{1}h_{3}+a_{j,k}(L_{2})b_{i,k}(L_{2})h_{2}h_{3}-a_{i,k}(L_{1})c_{j,k}(L_{2},p_{2})h_{1}h_{4}\\ +a_{j,k}(L_{2})c_{i,k}(L_{1},p_{1})h_{2}h_{4}-b_{i,k}(L_{1})c_{j,k}(L_{2},p_{2})h_{3}h_{4}-b_{j,k}(L_{2})c_{i,k}(L_{1},p_{1})h_{3}h_{4}\}\\ +\alpha_{ij}hd_{i}(L_{1})d_{j}(L_{2}),\\ K_{ij}^{2}=-\frac{\alpha_{ij}}{2}\sum_{k=1}^{K}\{a_{i,k}(L_{1})a_{j,k}(L_{2})h_{12}-b_{i,k}(L_{1})b_{j,k}(L_{2})h_{33}-c_{i,k}(L_{1},p_{1})c_{j,k}(L_{2},p_{2})h_{44}\\ -a_{i,k}(L_{1})b_{j,k}(L_{2})h_{13}+a_{j,k}(L_{2})b_{i,k}(L_{1})h_{23}\}+2\alpha_{ij}\nu_{ij}d_{i}(L_{1})d_{j}(L_{2}),$
and $\alpha_{ij}=\displaystyle\frac{\rho_{ij}\sigma_{i}\sigma_{j}}{2^{\nu_{ij}-1}\Gamma(\nu_{ij})}$.
@@ 3.2 Modelling vertical coefficients for the differential operator
where $\eta_{i,m}^{(k)}$ are coefficients and $B_{i,m}(p)$ are basis functions defined over the normalized depth domain $p\in[0,1]$. Following the construction in \@BBOPcite\@BAP\@BBNjun2008nonstationary\@BBCP and \@BBOPcite\@BAP\@BBNjun2011non\@BBCP, we consider Legendre polynomials as a natural choice for $B_{i,m}$. If $P_{m}(x)$ denotes the $m$th Legendre polynomial defined on $x\in[-1,1]$, we define $B_{i,m}(p)=P_{m}(2p-1)$ after rescaling the depth variable $p$ to the unit interval. This orthogonal basis provides smooth global representations well suited for broad vertical structures. As an alternative, B-spline basis functions $\mathcal{B}_{m}(p)$ may be used. B-splines offer localized control by joining piecewise polynomials at knots \@BBOPcitep\@BAP\@BBN(kunoth2018foundations)\@BBCP. Let $\mathcal{B}_{m}(p)$ denote the $m$th B-spline basis function of a chosen order. Then the basis expansion is written as $B_{i,m}(p)=\mathcal{B}_{m}(p)$. The placement and spacing of knots can be tailored to the variable of interest or the structure of the data, with closely spaced knots enabling finer resolution to capture sharp vertical transitions. However, using a large number of knots increases the number of parameters and can complicate estimation due to potential convergence issues.

### v4

@@ 1 Introduction
The core Argo data products are temperature and salinity–the two fundamental variables from which many other oceanographic variables can be derived, including freezing point, electrical conductivity, and viscosity \@BBOPcitep\@BAP\@BBN(pawlowicz2013key)\@BBCP, ocean heat content and potential density \@BBOPcitep\@BAP\@BBN(yarger2022functional)\@BBCP, and tropical cyclone-induced ocean thermal response \@BBOPcitep\@BAP\@BBN(hu2024spatiotemporal)\@BBCP. Temperature and salinity are central to understanding the oceans’ physical properties and dynamics: their distributions influence ocean circulation \@BBOPcitep\@BAP\@BBN(chen2022effects; gangopadhyay2022introduction)\@BBCP, shape climate processes \@BBOPcitep\@BAP\@BBN(olson2022effect)\@BBCP, and alter biogeochemistry \@BBOPcitep\@BAP\@BBN(galan2021argo; ding2022effects)\@BBCP. They also support diverse applications, including hydroacoustics \@BBOPcitep\@BAP\@BBN(jana2022sound)\@BBCP, offshore wind farm design \@BBOPcitep\@BAP\@BBN(escobar2016influence)\@BBCP, and assessments of environmental stressors in marine and freshwater systems \@BBOPcitep\@BAP\@BBN(walker2020warmer)\@BBCP.
@@ 1.1 Spatial covariance models in 2D and 3D space
$C_{ij}(\mathbf{s}_{1},\mathbf{s}_{2})=\rho_{ij}\frac{\sigma_{i}(\mathbf{s}_{1})\sigma_{j}(\mathbf{s}_{2})}{|\mathbf{D}_{ij}(\mathbf{s}_{1},\mathbf{s}_{2})|^{1/2}}\mathcal{M}_{\nu_{ij}}\left\{\left((\mathbf{s}_{1}-\mathbf{s}_{2})^{\top}\mathbf{D}_{ij}(\mathbf{s}_{1},\mathbf{s}_{2})^{-1}(\mathbf{s}_{1}-\mathbf{s}_{2})\right)^{1/2}\right\},$ (5)
where $\mathbf{D}_{ij}(\mathbf{s}_{1},\mathbf{s}_{2})=\frac{1}{2}\left\{\mathbf{D}_{i}(\mathbf{s}_{1})+\mathbf{D}_{j}(\mathbf{s}_{2})\right\}$ is a $d\times d$ positive definite kernel matrix that controls spatially varying anisotropy, for $i,j=1,\ldots,q$. Despite this innovation, extending the Matérn model to include nonstationary colocated correlations $\rho_{ij}(\mathbf{s})$ remains challenging. One proposed remedy is to replace the constants $\sigma_{i}^{2}$ and $\beta_{ij}$ with kernel-smoothed nonparametric estimators \@BBOPcitep\@BAP\@BBN(kleiber2012nonstationary)\@BBCP:
@@ 2.1 Spatial Interpolation of Argo Data
One widely used Argo product is the Roemmich–Gilson Argo climatology from the Scripps Institution of Oceanography \@BBOPcitep\@BAP\@BBN(roemmich2009)\@BBCP, which employs weighted local regression. Let $\mathbf{s}=(L,l,p)$ denote spatial location, where $L$ is latitude, $l$ is longitude, and $p$ is pressure (commonly treated as depth in the upper ocean). Let $t$ denote time in yeardays (days since January 1 of a given year). For a reference location $\mathbf{s}_{0}=(L_{0},l_{0},p_{0})^{\top}$, \@BBOPcite\@BAP\@BBNroemmich2009\@BBCP fit the mean function:
$\mu_{\mathbf{s}_{0}}(\mathbf{s},t)=\beta_{0}+\beta_{1}(L-L_{0})+\beta_{2}(l-l_{0})+\beta_{3}(L-L_{0})^{2}+\beta_{4}(l-l_{0})^{2}+\beta_{5}(p-p_{0})\\ +\beta_{6}(p-p_{0})^{2}+\sum_{k=1}^{6}\gamma_{k}\sin\left(2\pi k\frac{t}{365.25}\right)+\sum_{k=1}^{6}\delta_{k}\cos\left(2\pi k\frac{t}{365.25}\right)$ (7)
using observations from the nearest horizontal neighbors, pressure levels and each of the 12 calendar months, drawn from the entire Argo record. Here, $\beta_{0}$ and $\beta_{k}$, $\gamma_{k}$, and $\delta_{k}$, $k=1,\ldots,6$, are scalar coefficients. The residuals were modeled using the spatial covariance function:
@@
Most studies discretize the vertical domain into fixed pressure levels and interpolate each layer independently, which requires preprocessing irregular profile data and can introduce systematic error \@BBOPcitep\@BAP\@BBN(kuusela2018locally; hu2024spatiotemporal)\@BBCP. To avoid this, \@BBOPcite\@BAP\@BBNyarger2022functional\@BBCP treated Argo profiles as functions of pressure and modeled the local mean as
@@ 3.1 Multivariate Nonstationary Spatial Models in 3D Domain
Furthermore, suppose the cross-covariance function $C_{ij}^{X}$ is chosen to be the parsimonious Matérn cross-covariance function in (1) with smoothness $\nu_{ij}+1$. Then, the multivariate process $\mathbf{Z}=(Z_{1},Z_{2},\ldots,Z_{q})^{\top}$ has a nonstationary cross-covariance function:
@@
$K_{ij}^{1}=\frac{\alpha_{ij}}{4}\sum_{k=1}^{K}\{a_{i,k}(L_{1})a_{j,k}(L_{2})h_{1}h_{2}-b_{i,k}(L_{1})b_{j,k}(L_{2})h_{3}^{2}-c_{i,k}(L_{1},p_{1})c_{j,k}(L_{2},p_{2})h_{4}^{2}\\ -a_{i,k}(L_{1})b_{j,k}(L_{2})h_{1}h_{3}+a_{j,k}(L_{2})b_{i,k}(L_{1})h_{2}h_{3}-a_{i,k}(L_{1})c_{j,k}(L_{2},p_{2})h_{1}h_{4}\\ +a_{j,k}(L_{2})c_{i,k}(L_{1},p_{1})h_{2}h_{4}-b_{i,k}(L_{1})c_{j,k}(L_{2},p_{2})h_{3}h_{4}-b_{j,k}(L_{2})c_{i,k}(L_{1},p_{1})h_{3}h_{4}\}\\ +\alpha_{ij}hd_{i}(L_{1})d_{j}(L_{2}),\\ K_{ij}^{2}=-\frac{\alpha_{ij}}{2}\sum_{k=1}^{K}\{a_{i,k}(L_{1})a_{j,k}(L_{2})h_{12}-b_{i,k}(L_{1})b_{j,k}(L_{2})h_{33}-c_{i,k}(L_{1},p_{1})c_{j,k}(L_{2},p_{2})h_{44}\\ -a_{i,k}(L_{1})b_{j,k}(L_{2})h_{13}+a_{j,k}(L_{2})b_{i,k}(L_{1})h_{23}\}+2\alpha_{ij}\nu_{ij}d_{i}(L_{1})d_{j}(L_{2}),$
and $\alpha_{ij}=\displaystyle\frac{\rho_{ij}\sigma_{i}\sigma_{j}}{2^{\nu_{ij}}\Gamma(\nu_{ij}+1)}$.
@@ 3.2 Modelling vertical coefficients for the differential operator
where $\eta_{i,m}^{(k)}$ are coefficients and $B_{i,m}(p)$ are basis functions defined over the normalized depth domain $p\in[0,1]$. Following the construction in \@BBOPcite\@BAP\@BBNjun2008nonstationary\@BBCP and \@BBOPcite\@BAP\@BBNjun2011non\@BBCP, we consider Legendre polynomials as a natural choice for $B_{i,m}$. If $P_{m}(x)$ denotes the $m$th Legendre polynomial defined on $x\in[-1,1]$, we define $B_{i,m}(p)=P_{m}(2p-1)$ after rescaling the depth variable $p$ to the unit interval. This orthogonal basis provides smooth global representations well suited for broad vertical structures. As an alternative, B-spline basis functions $\mathcal{B}_{m}(p)$ may be used. B-splines offer localized control by joining piecewise polynomials at knots \@BBOPcitep\@BAP\@BBN(kunoth2018splines)\@BBCP. Let $\mathcal{B}_{m}(p)$ denote the $m$th B-spline basis function of a chosen order. Then the basis expansion is written as $B_{i,m}(p)=\mathcal{B}_{m}(p)$. The placement and spacing of knots can be tailored to the variable of interest or the structure of the data, with closely spaced knots enabling finer resolution to capture sharp vertical transitions. However, using a large number of knots increases the number of parameters and can complicate estimation due to potential convergence issues.